### v2

# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

Abstract
Optical tweezer arrays have had a transformative impact on atomic and molecular physics over the past years, and they now form the backbone for a wide range of leading experiments in quantum computing, simulation, and metrology. Underlying this development is the simplicity of single particle control and detection inherent to the technique. Typical experiments trap tens to hundreds of atomic qubits, and very recently systems with around one thousand atoms were realized without defining qubits or demonstrating coherent control. However, scaling to thousands of atomic qubits with long coherence times and low-loss, high-fidelity imaging is an outstanding challenge and critical for progress in quantum computing, simulation, and metrology, in particular, towards applications with quantum error correction. Here, we experimentally realize an array of optical tweezers trapping over 6,100 neutral atoms in around 12,000 sites while simultaneously surpassing state-of-the-art performance for several key metrics associated with fundamental limitations of the platform. Specifically, while scaling to such a large number of atoms, we also demonstrate a coherence time of 12.6(1) seconds, a record for hyperfine qubits in an optical tweezer array. Further, we show trapping lifetimes close to 23 minutes in a room-temperature apparatus, enabling record-high imaging survival of $99.98952(1)\%$ in combination with an imaging fidelity of over $99.99\%$. Our results, together with other recent developments, indicate that universal quantum computing with ten thousand atomic qubits could be a near-term prospect. Furthermore, our work could pave the way for quantum simulation and metrology experiments with inherent single particle readout and positioning capabilities at a similar scale.
Optical tweezer arrays browaeys_many-body_2020, kaufman_quantum_2021 have transformed atomic and molecular physics experiments by simplifying detection and enabling control at the individual-particle level barredo_atom-by-atom_2016, endres_atom-by-atom_2016, kim_situ_2016, liu_building_2018, anderegg_optical_2019, resulting in rapid, recent progress in quantum computing saffman_quantum_2016, henriet_quantum_2020, bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024, quantum simulation bernien_probing_2017, browaeys_many-body_2020, ebadi_quantum_2021, scholl_quantum_2021, shaw_benchmarking_2024, and optical frequency metrology norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020. They are part of broader efforts towards single atom control in optical traps, including lattices miroshnychenko_atom-sorting_2006, weitenberg_single-spin_2011 and lattice-tweezer hybrid systems young_half-minute-scale_2020, tao_high-fidelity_2023, gyger_continuous_2024. For quantum computers, simulators of spin systems and tweezer clocks, each atom typically encodes a single qubit that is controlled with electromagnetic fields, and ideally features long coherence times to enable these applications with high fidelity. Such optically trapped atomic qubits coexist with other platforms that have single qubit control and readout, including ion traps bruzewicz_trapped-ion_2019 and superconducting qubits kjaergaard_superconducting_2020.
There are important incentives to scale up such fully programmable qubit platforms. For example, optical clocks gain in stability with increasing atom number ludlow_optical_2015, rosenband_exponential_2013, and proposals have been made for quantum simulation experiments that benefit from several thousand qubits to explore emergent collective behavior orourke_entanglement_2023, julia-farre_amorphous_2024 or to demonstrate verifiable quantum advantage haferkamp_closing_2020, ringbauer_verifiable_2023.
In particular, quantum error correction (QEC) requires large system sizes, while maintaining high fidelities and long coherence times, in order to scale up the number of logical qubits preskill_quantum_2018, google_quantum_ai_suppressing_2023. Even the most resource-efficient QEC protocols require at least several thousand physical qubits with an error-per-gate level of $10^{-3}$ to encode $>$100 error-corrected logical qubits bravyi_high-threshold_2023, xu_constant-overhead_2023. However, published results from universal quantum computing architectures based on ion traps and superconducting qubits currently leverage only tens to hundreds of qubits moses_race_2023, kim_evidence_2023, and with current technologies for fabrication and control, generally suffer from deleterious effects as they are scaled kjaergaard_superconducting_2020, bruzewicz_trapped-ion_2019.
Neutral atoms in optical tweezer arrays are a promising platform towards rapid scalability in the near-term thanks to a programmable architecture that is readily adaptable to larger system sizes. In particular, universal quantum computing has recently been realized in such systems bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024, based on demonstrations of individual qubit addressing huie_repetitive_2023, lis_midcircuit_2023, norcia_midcircuit_2023, shaw_multi-ensemble_2024, high-fidelity entangling gates evered_high-fidelity_2023, finkelstein_universal_2024, and coherence-preserving dynamical reconfigurability beugnon_two-dimensional_2007, bluvstein_quantum_2022, finkelstein_universal_2024, alongside ancilla-based mid-circuit measurement singh_mid-circuit_2023, bluvstein_logical_2024, finkelstein_universal_2024 and erasure error detection scholl_erasure_2023, ma_high-fidelity_2023. Leveraging the versatility of the optical tweezer architecture, several applications have been shown, including executing quantum phase estimation graham_multi-qubit_2022, realizing the toric code structure bluvstein_quantum_2022, building an optical clock with ancilla-based readout finkelstein_universal_2024, and implementing logical qubit operations bluvstein_logical_2024.
In terms of current system sizes, tens to hundreds of qubits are controlled by typical tweezer array experiments young_half-minute-scale_2020, scholl_quantum_2021, huft_simple_2022, singh_mid-circuit_2023, tao_high-fidelity_2023, bluvstein_logical_2024, shaw_benchmarking_2024. Very recently, tweezer systems with about a thousand atoms have been realized in a discontiguous array based on interleaved microlens elements pause_supercharged_2024, and via repeated reloading from a reservoir norcia_iterative_2024, gyger_continuous_2024, following earlier iterative reloading schemes singh_dual-element_2022, shaw_dark-state_2023; none of these experiments, however, report control of qubits or measurement of coherence times.
Figure 1: Large-scale tweezer array. a, Representative single-shot image of single cesium atoms across a 12,001 tweezer array. Inset: magnified view of a subsection of the stochastically loaded array. b, Averaged image of single atoms across a 12,001 tweezer array. Inset: magnified view of a subsection of the averaged array. Atoms are spaced by ${\sim}7.15\,\mu$m and held in 1061 nm and 1055 nm optical tweezers. c, Schematic of the optical tweezer array generation. Tweezer arrays, generated by two spatial light modulators (SLM), at 1061 nm and 1055 nm are combined with orthogonal polarization, and focused through an objective with a numerical aperture (NA) of 0.65 and a field of view (FOV) of 1.5 mm in diameter. The $y$-axis in the figure corresponds to the direction of gravity. We collect scattered photons from single atoms through the same objective (NA$=0.55$ at the imaging wavelength) and image them on a qCMOS camera. d, Histogram of filling fraction. We load 6,139 single atoms on average per experimental iteration. This corresponds to loading 51.2% of the array on average, with a relative standard deviation of 1.13% over 16,000 iterations. e, Summary of the key metrics demonstrated in this work. The imaging survival quoted is a steady-state imaging survival probability, and the global single-qubit gate fidelity is obtained from global randomized benchmarking (RB), as detailed in the text.
Here, we demonstrate a tweezer array with around 12,000 sites, that traps over 6,100 atomic qubits, simultaneously surpassing state-of-the-art values for metrics associated with fundamental limitations of the platform, including hyperfine qubit coherence time and trapping lifetime in a room-temperature apparatus, as well as combined imaging fidelity and survival (Fig. 1). Our results have implications for aforementioned applications in quantum science, in particular, concerning large-scale quantum computing and error correction, as discussed in more detail below.
Summary of approach and results
Our approach leverages high-power trapping of single atoms at far-off-resonant wavelengths in a specially designed, room-temperature vacuum chamber (Methods, Ext. Data Fig. 1a), enabling low-loss, high-fidelity imaging in combination with long hyperfine coherence times at the scale of 6,100 qubits (Fig. 1e). Specifically, we demonstrate imaging of single cesium-133 atoms with a survival probability of $99.98952(1)\%$, in combination with an imaging fidelity of $99.99374(8)\%$, surpassing the state-of-the-art achieved in much smaller arrayscovey_2000-times_2019. This, alongside a 22.9(1) minute vacuum-limited lifetime in our room-temperature apparatus schymik_single_2021 — over three times longer than previous records for tweezer arrays in room-temperature apparatuses covey_2000-times_2019 — provides realistic timescales for array operations in large scale arrays with minimal loss, e.g., for atomic rearrangement barredo_atom-by-atom_2016, endres_atom-by-atom_2016, kim_situ_2016.
Importantly, we further demonstrate a coherence time of 12.6(1) s, a record for a hyperfine qubit tweezer array, surpassing previous values by almost an order of magnitude bluvstein_quantum_2022, graham_multi-qubit_2022 and approaching results for a single hyperfine qubit in a customized blue-detuned trap tian_coherence_2023, alkali atoms in an optical lattice dudin_light_2013, and nuclear qubits in a tweezer array barnes_assembly_2022. We also show a single-qubit gate fidelity of $99.9834(2)\%$ measured with global randomized benchmarking, limited by technical noise. We observe high uniformity across the sites in the array for loading probability (Ext. Data Fig. 2c), trap depth (Ext. Data Fig. 2d), imaging survival (Ext. Data Fig. 4a), scattering rate (Ext. Data Fig. 3d), and qubit frequency (Ext. Data Fig. 5d). Our results, when paired with other recent techniques for scaling single-site addressing zhang_scaled_2024, along with demonstrations of two-qubit gate operations evered_high-fidelity_2023, ma_high-fidelity_2023, finkelstein_universal_2024 and qubit rearrangement beugnon_two-dimensional_2007, bluvstein_quantum_2022, indicate that high-fidelity universal quantum computing with about ten thousand atomic qubits could be a near-term prospect, providing a path towards QEC with hundreds of logical qubits xu_constant-overhead_2023.
Large-scale optical tweezer generation
We scale the optical tweezer array platform to around 12,000 tweezers, while prioritizing achieving long qubit coherence times and high fidelities (Fig. 1e). As such, we trap using near-infrared wavelengths, far-detuned from dominant electric-dipole transitions to minimize hyperfine qubit decoherence from photon scattering and dephasing processes ozeri_hyperfine_2005, ozeri_errors_2007. Cesium atoms possess the highest polarizability among the stable alkali metal atoms at near-infrared wavelengths where commercial fiber amplifiers routinely provide continuous-wave laser powers that exceed 100 W. Thus, a large number of traps can be generated with sufficient depth. A representative single shot image of the array is shown in Fig. 1a, and an averaged image is shown in Fig. 1b.
The atoms are spaced by ${\sim}7.15\,\mathrm{\mu m}$ and held in traps at 1055 nm and 1061 nm, generated using spatial light modulators (SLMs), whose hologram phases are optimized with a weighted Gerchberg-Saxton algorithm kim_gerchberg-saxton_2019, kim_large-scale_2019 to uniformize the tweezer trap depth across the array (Methods). The tweezer light is combined with polarization and focused through a high numerical aperture objective with an unusually large field of view of 1.5 mm diameter that provides a large area for qubit trapping and manipulation in ultra-high vacuum (Fig. 1c). We measure an average trap depth of $k_{B}\times 0.20(1)$ mK, with a standard deviation of $11.4\%$ across all sites (Ext. Data Fig. 2d), enabling consistent loading probability per site.
Loading and imaging single atoms
We demonstrate uniform loading and high imaging fidelity across the sites in the array. To load single atoms in the tweezers, we cool and then parity-project schlosser_sub-poissonian_2001 from a ${\sim}1.6$ mm $1/e^{2}$ diameter magneto-optical trap. Before imaging the atoms, we use a multi-pronged approach to filter out atoms in spurious off-plane traps, residual from the SLM tweezer creation (Methods).
Figure 2: High-fidelity atom detection in a large-scale tweezer array. Imaging histogram showing the number of photons collected per site and per image. Note that the horizontal axes are weighted photon counts (see text); for non-weighted photon counts, see Ext. Data Fig. 3b. a, Imaging histogram of three randomly selected sites in the array (where $x$ and $y$ respectively denote the horizontal and vertical site indices in the array), and b, averaged over all sites in the array. Per-site histograms are fitted with a Poissonian model that integrates losses during imaging (Methods). The wide separation of peaks for empty and filled tweezers enables the high imaging fidelity presented in this work. The binarization threshold used to determine tweezer occupation is indicated by the vertical dashed line, and the average point-spread functions for the two classifications (atom absent and atom detected) are shown next to their corresponding peaks. Inset: the same histogram presented with a log-scale vertical axis. The weighted average relative error bar per bin is 0.08% (0.05% for the log-scale inset due to the smaller number of bins). Data for the histogram is collected from 16,000 iterations of the experiment.
We then zero the magnetic field and apply two-dimensional polarization gradient cooling (2D PGC) beams that are parallel to the atom array plane ($x$-$z$ plane in Fig. 1c) for fluorescence imaging of single atoms, which also simultaneously cool the atoms. The photons scattered from each atom are collected on a quantitative CMOS (qCMOS) camera for 80 ms, using the same objective with which the traps are created. We find that each site is loaded with a single atom on average 51.2% of the time with a relative standard deviation of 3.4% across the sites, demonstrating uniform filling of single atoms in a large tweezer array (Ext. Data Fig. 2c). This allows us to load over 6,100 sites on average in each iteration (Fig. 1d).
We detect and distinguish atomic presence in the array with high fidelity. Each image undergoes a binarization procedure (detailed in Methods) whereby each site is attributed a value of 0 (no atom detected) or 1 (one atom detected). Note that we detect no more than one atom schlosser_sub-poissonian_2001 in each tweezer. We weight the collected photons in a $7\times 7$-pixel box centered around each site cooper_alkaline-earth_2018, so as to add more weight to pixels close to the center of each site’s point-spread function (Ext. Data Fig. 3a). The resulting signal is compared with a threshold to determine if an atom is present or not (Fig. 2).
We characterize the imaging fidelity, defined as the probability of correctly labeling atomic presence in a site, with a model-free approach, where no assumption is made about the photon distribution from Fig. 2. To this end, we identify anomalous series of binary outputs norcia_microscopic_2018 in three consecutive images. For instance, $0\rightarrow 1\rightarrow 1$ would point to a false negative event in the first image, while $1\rightarrow 1\rightarrow 0$ could be due to atom loss during the second image or a false negative event in the third one. This approach allows us to precisely decouple inherent atom loss from false negatives or positives (Methods). From this we find an imaging fidelity of 99.99374(8)%. Crucial to this result are the homogeneous photon scattering rate across the array (Ext. Data Fig. 3d) and the consistency of the point-spread function even for sites at the edge of the array (waist radius of 1.7 pixels with a standard deviation of 0.2 pixels). Furthermore, we find that treating each site separately with an individual threshold only marginally improves the imaging fidelity to 99.9939(1)%, indicating that the imaging parameters are sufficiently consistent across the atom array. Finally, we estimate that the imaging fidelity in the absence of atomic loss would be closer to 99.999% (Methods).
Imaging survival and vacuum-limited lifetime
The probability of not losing a single atom in a tweezer array during imaging and due to finite vacuum lifetime both decrease exponentially in the number of atoms in an array, making these crucial metrics to optimize for large-scale array operation. The vacuum-limited lifetime, in particular, quantifies a fundamental experimental limitation on the amount of time in which operations can be executed without loss of an atom. This can, for example, be applied to give an upper limit on the fidelity with which one can achieve a defect-free array via atom rearrangement barredo_atom-by-atom_2016, endres_atom-by-atom_2016.
We probe the vacuum-limited lifetime by using an empirically optimized cooling sequence consisting of a 10 ms 2D PGC cooling block every 2 s. By fitting the exponential decay of the atom survival, we find a $1/e$ lifetime of 22.9(1) min (Fig. 3a). This is more than three times longer than the longest reported lifetime for neutral atoms in an optical tweezer array using a room-temperature apparatus covey_2000-times_2019, and only within a factor of 5 of the longest reported lifetime in a cryogenic apparatus schymik_single_2021. The result indicates that the probability of losing a single atom across the entire array remains under 50% during 100 ms, a relevant timescale for dynamical array reconfiguration and quantum processor operation.
Figure 3: Long vacuum-limited lifetime and high imaging survival probability a, Vacuum-limited lifetime. Array-averaged survival fraction as a function of hold time is plotted. Two experiments are shown in the figure: with and without cooling. The green markers show data with a 10 ms 2D PGC block applied every 2 s (pulsed PGC), and the blue markers show the data without cooling (hold). The error bars are smaller than the markers. We find a $1/e$ lifetime of around 2.2 min without cooling. When the pulsed PGC block is applied, by fitting the data with $p(t)\propto\exp(-t/\tau)$, we find a vacuum lifetime of $\tau=22.9(1)$ min. b, Array-averaged survival fraction after many successive images. Between each image, we hold the atoms for 10 ms, without applying any cooling beams. We fit the data with $p(N)\propto p_{1}^{N}$, where $p(N)$ is the survival fraction after imaging $N$ times. From the fit, we find a steady-state imaging survival probability of $p_{1}=0.9998952(1)$. The light purple fill shows the estimated 67% confidence interval.
Figure 4: Long coherence times and high-fidelity single-qubit gates in a large atom array. a, Array-averaged Rabi oscillations between the hyperfine clock states $|0\rangle$ and $|1\rangle$. The fitted Rabi frequency is 24.611(1) kHz. The observed decay after several hundred microseconds stems from the spatially-varying Rabi frequency (Ext. Data Fig. 5c). b, Array-averaged Ramsey oscillations. During free evolution, the microwave drive field is detuned by 1 kHz, resulting in Ramsey oscillations. The characteristic decay time of these oscillations is $T_{2}^{*}=14.0(1)\,\mathrm{ms}$ fitting the average signal of all atoms. The light blue dashed line shows the decay time $T_{2}^{*(\mathrm{site})}=25.5\ \mathrm{ms}$ from fitting individual sites first and averaging the decay time afterwards. c, Measurement of the dephasing time $T_{2}$ after dynamical decoupling. After an initial $\pi/2$ pulse, a variable number of XY16 dynamical decoupling cycles with a fixed time $\tau=12.5\ \mathrm{ms}$ between $\pi$ pulses are used to offset the reversible dephasing. The phase of the final $\pi/2$ pulse is chosen to be either 0 or $\pi$, and subtracting the population difference in these two cases provides the coherence contrast. The contrast decay is fitted to obtain $T_{2}=12.6(1)\,\mathrm{s}$. d, Randomized benchmarking of the global single-qubit gate fidelity. For each number of Clifford gates, 60 different random gate strings of this length are applied, after which the overall inverse of the string is applied. The translucent markers represent the return probability for each string of gates, while the solid green markers indicate the averaged return probability over the 60 different strings. The decay of the final population to $1/2$ is fitted to $(1-d)^{N}$ and $F_{c}=1-d/2$ represents the average Clifford gate fidelity.
Moreover, we accurately characterize the imaging survival probability, without assuming any parameters, by performing 80 ms repeated imaging up to 1000 times, after which ${\sim}90$% of initially loaded atoms still survive (Fig. 3b). This corresponds to a steady-state imaging survival probability of $99.98952(1)\%$, mostly limited by vacuum lifetime. This, to the best of our knowledge, surpasses prior studies reporting record steady-state imaging survival using single alkaline-earth metal covey_2000-times_2019 and alkali-metal blodgett_imaging_parsed_2023 atoms in optical tweezers. These results, and the uniformity of imaging survival across the array (Ext. Data Fig. 4a), enable low-loss, high-fidelity detection of single atoms in large-scale arrays, crucial components for the practical use of the system (see Discussion and Outlook).
Qubit coherence
Key to recent progress in quantum computing and metrology with neutral atoms is the ability to encode a qubit in long-lived states of an atom, such as hyperfine states bluvstein_quantum_2022, graham_multi-qubit_2022, nuclear spin states barnes_assembly_2022, huie_repetitive_2023, lis_midcircuit_2023, norcia_midcircuit_2023, or optical clock states norcia_seconds-scale_2019, madjarov_atomic-array_2019. In cesium atoms, the hyperfine ground states ($|F=3,\ m_{F}=0\rangle\equiv|0\rangle$ and $|F=4,\ m_{F}=0\rangle\equiv|1\rangle$) provide such a subspace for storing quantum information. Furthermore, entanglement via Rydberg interactions can be readily transferred to this qubit to realize high-fidelity two-qubit gates evered_high-fidelity_2023, finkelstein_universal_2024. We now demonstrate the storage and manipulation of quantum information in a large-scale atom array by measuring the coherence time and global single-qubit gate fidelity. To this end we use a microwave horn to drive the hyperfine transition (Fig. 4a).
Preserving the coherence of a quantum system as it is scaled up is a known challenge across platforms for quantum computing and simulation kjaergaard_superconducting_2020. This difficulty persists even for neutral atoms, albeit at a lower level, due to residual interactions with a noisy and inhomogeneous electromagnetic environment, particularly with tweezers themselves (due to induced scattering and dephasing). Using far-off-resonant optical tweezers helps to preserve coherence, since the differential light shift of the hyperfine qubit decreases as $1/\Delta_{\mathrm{tweezer}}$ at constant trap depth, and the scattering rate as $1/\Delta_{\mathrm{tweezer}}^{3}$, where $\Delta_{\mathrm{tweezer}}$ is the tweezer laser detuning relative to the dominant electronic transition kuhr_analysis_2005, ozeri_hyperfine_2005, ozeri_errors_2007. We verify this by measuring a depolarization time of $T_{1}=119(1)$ s (Ext. Data Fig. 5e), and an ensemble dephasing time in the atom array of $T_{2}^{*}=14.0(1)$ ms (Fig. 4b), limited by trap depth inhomogeneity. Measured site-by-site, the dephasing time is $T_{2}^{*(\mathrm{site})}=25.5$ ms, consistent with being limited by an atomic temperature of ${\sim}4.8\ \mathrm{\mu K}$ during microwave operation kuhr_analysis_2005. We also measure the standard deviation of the qubit frequency in the array to be $1.5\times 10^{-9}$, or 14 Hz in absolute value (Ext. Data Fig. 5d).
The remaining reversible dephasing can be further mitigated by dynamical decoupling. By applying cycles of XY16 sequences gullion_new_1990, souza_robust_2012 with a dwell time of 12.5 ms between $\pi$ pulses, the measured dephasing time is $T_{2}=12.6(1)$ s, thereby setting a new benchmark for the coherence time of an array of hyperfine qubits in optical tweezers bluvstein_quantum_2022, graham_multi-qubit_2022 (Fig. 4c). Combined with recent advances in neutral atom quantum processors based on qubit transport beugnon_two-dimensional_2007, bluvstein_quantum_2022, bluvstein_logical_2024, which require dozens of microseconds for a set of parallel gates (and thus long coherence times), such long coherence times could enable the realization of complex quantum algorithms with thousands of qubits in the near-term.
Finally, we demonstrate global coherent single-qubit operation of our atom array and probe single-qubit gate fidelities through global randomized benchmarking knill_randomized_2008, xia_randomized_2015, nikolov_randomized_2023. To compensate for the inhomogeneous Rabi frequency across the array, which would result in pulse-length errors, we use the SCROFULOUS composite pulse cummins_tackling_2003. We apply gates sampled from the 24 unitaries composing the Clifford group $C_{1}$, followed by an inverse operation, and measure the final population in $|1\rangle$ (Fig. 4d). Fitting the decay as the number of gates increases yields an average Clifford gate fidelity $F_{c}=0.999834(2)$, limited by phase noise in our system likely due to magnetic field noise (Methods). This could be readily addressed by upgrading the current sources driving the magnetic field coils, shielding the vacuum cell, or by operating at MHz-scale by driving optical Raman transitions levine_dispersive_2022.
Discussion and outlook
We have demonstrated scaling of the optical tweezer array platform to over 6,100 individually-trapped atoms. While some crucial ingredients have yet to be demonstrated, especially the ability to rearrange atoms and entangle atomic qubits, we show equal or superior results to state-of-the-art experiments on several key metrics that are known bottlenecks for quantum computing, quantum simulation and quantum metrology. Notably, we achieve record-high imaging survival covey_2000-times_2019 alongside competitive imaging fidelity. We measure record room-temperature vacuum-limited lifetime covey_2000-times_2019 and coherence time in alkali metal atom tweezer arrays bluvstein_quantum_2022, graham_multi-qubit_2022, with a high global single-qubit gate fidelity, only limited by technical noise. We thereby largely circumvent aspects of degradation in classical and quantum control commonly faced by platforms as they are scaled up.
Our results usher in a new generation of neutral atom quantum processors based on several thousand of qubits, as well as large-scale programmable devices enabling advances in quantum metrology rosenband_exponential_2013, ludlow_optical_2015, norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020, finkelstein_universal_2024 and simulation haferkamp_closing_2020, orourke_entanglement_2023, julia-farre_amorphous_2024. For example, our platform — with the demonstrated qubit numbers — could be used for verifiable quantum advantage with low-depth evolution haferkamp_closing_2020, ringbauer_verifiable_2023. Tweezer clocks could be scaled using near-infrared, high-power tweezers for loading and imaging tao_high-fidelity_2023 before transferring atoms to magic-wavelength traps for clock operation norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020, finkelstein_universal_2024. We also foresee applications in quantum simulation for problems where boundary effects play an important role bernien_probing_2017, browaeys_many-body_2020, ebadi_quantum_2021, scholl_quantum_2021, orourke_entanglement_2023, which can be minimized with the large system sizes demonstrated here. In order to realize directions requiring interactions between atoms, we plan on exciting to Rydberg states saffman_quantum_2010. Promisingly, cesium provides a two-photon excitation pathway through the $7\textrm{P}_{3/2}$ intermediate state with a 455 nm lower photon transition and 1060 nm upper photon transition. With a large detuning from the intermediate state to minimize scattering, and with widely available laser powers at these wavelengths, it is possible to achieve Rabi frequencies as high as 5 MHz with minimal inhomogeneity across the full field of view.
Concerning universal quantum computation, the unusually large 1.5 mm field of view (FOV) of our objective provides enough space to engineer zones around our current array (900 $\mathrm{\mu m}$ in diameter) for interaction, storage, qubit rotation, and mid-circuit readout, enabling operations on qubits without disturbing adjacent atoms beugnon_two-dimensional_2007, bluvstein_quantum_2022. In order to utilize such a zone architecture for quantum computation, it is imperative to dynamically rearrange atoms across the optical tweezer array barredo_atom-by-atom_2016, endres_atom-by-atom_2016. Using commercially available acousto-optic deflectors with large field-of-view (${>}500\,\mathrm{\mu m}$ with our optical parameters) and fast steering mirrors, we foresee a straightforward pathway to low-loss transport of atoms across the complete objective field-of-view of 1.5 mm. This could be combined with recent technical developments that allow for high-speed parallel addressing of 10,000 sites zhang_scaled_2024, and demonstrations of ancilla-based mid-circuit readout techniques singh_mid-circuit_2023, bluvstein_logical_2024, finkelstein_universal_2024, to ultimately pave the way for repeated quantum error correction in neutral atom arrays with thousands of physical qubits.
For most quantum simulation and computation tasks, rearrangement of the stochastically loaded atom array into a defect-free array is required barredo_atom-by-atom_2016, endres_atom-by-atom_2016. The probability of obtaining a defect-free array can be limited by several factors, including finite imaging fidelity, imaging survival, imperfections in the rearrangement sequence, or finite vacuum-limited lifetime. We can provide an early estimate assuming a rearrangement time much less than our vacuum lifetime, that based on imaging survival alone, the probability to obtain a defect-free array of  6,040 rearranged atoms can be up to 39.5%, with optimal choices of imaging thresholding for this purpose (Ext. Data Fig. 4b). The average total number of defects corresponding to this estimation is 0.93.
Finally, our work indicates that further scaling of the optical tweezer array platform to tens of thousands of trapped atoms should be achievable with current technology, while essentially preserving high-fidelity control. In our present apparatus, several factors limit the number of sites. One limitation is the finite number of pixels of each SLM (which reduces the diffraction efficiency as the array size is increased), along with reduced SLM diffraction efficiency at higher incident laser powers. By using available higher-resolution SLMs, and by exploring techniques with higher pixel modulation depth moreno_diffraction_2020, we hope to utilize both power and field of view more efficiently. Furthermore, we observe worsening optical aberrations at tweezer powers greater than that in the present study due to thermal heating of the objective. This is the main limitation on atom number for the results in this work, even after aberrations were mitigated using the SLM (Methods). This constraint could be circumvented in the short term by dissipating heat from the objective more effectively to suppress detrimental aberrations and in the long term by utilizing an objective with a housing material that retains less heat; a combination of these upgrades should allow us to almost double the number of tweezers that we create using two fiber amplifiers. We further anticipate the potential to switch from polarization combination to wavelength-based array combination, opening up further avenues for increasing tweezer number with similar techniques. Atom numbers may further be increased in our array with the same number of tweezers by utilizing enhanced loading brown_gray-molasses_2019 or re-loading techniques shaw_dark-state_2023. Already in the near-term, we expect to increase the number of atomic qubits to over ten thousand with the current system using a subset of these techniques.
Acknowledgements.
We acknowledge insightful discussions with, and feedback from, Adam Shaw, Ran Finkelstein, Pascal Scholl, Joonhee Choi, and Soonwon Choi. We acknowledge support from the Gordon and Betty Moore Foundation (Grant GBMF11562), the Weston Havens Foundation, the Institute for Quantum Information and Matter, an NSF Physics Frontiers Center (NSF Grant PHY-1733907), the NSF QLCI program (2016245), the NSF CAREER award (1753386), the Army Research Office MURI program (W911NF2010136), the U.S. Department of Energy (DE-SC0021951), the DARPA ONISQ program (W911NF2010021), and the Air Force Office for Scientific Research Young Investigator Program (FA9550-19-1-0044). Support is also acknowledged from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator. HJM acknowledges support from the NSF Graduate Research Fellowship Program under Grant No. 2139433. KHL acknowledges support from the AWS-Quantum postdoctoral fellowship and the NUS Development Grant AY2023/2024.
Data availability
The data and codes that support the findings of this study are available from the corresponding author upon request.